In a two levels system, the levels are separated by an energy E2 - E1. Find the ratio of population inversions at room temperature if transition between levels occur at (i) λ = 0.55 μm (ii) λ = 3 μm


1
Expert's answer
2021-04-12T07:03:51-0400

i)


"\frac{A_{21}}{B_{12}}=\frac{2h}{\lambda^3}=\frac{2(6.6\cdot10^{-34})}{(0.55\cdot10^{-6})^3}=7.9\cdot10^{-15}\frac{Js}{m^3}"

2)


"\frac{A_{21}}{B_{12}}=\frac{2h}{\lambda^3}=\frac{2(6.6\cdot10^{-34})}{(3\cdot10^{-6})^3}=4.9\cdot10^{-17}\frac{Js}{m^3}"